# The Coordination Gap: Alternation Metrics for Temporal Dynamics in Multi-Agent Battle of the Exes

### 06 Mar 2026 v1

@@ Abstract
Multi-agent coordination dilemmas expose a fundamental tension between individual optimization and collective welfare, yet characterizing such coordination requires metrics sensitive to temporal structure and collective dynamics. As a diagnostic testbed, we study a BoE-derived multi-agent variant of the Battle of the Exes, formalizing it as a Markov game in which turn-taking emerges as a periodic coordination regime. Conventional outcome-based metrics (e.g., efficiency and min/max fairness) are temporally blind—they cannot distinguish structured alternation from monopolistic or random access patterns—and fairness ratios lose discriminative power as $n$ grows, obscuring inequities.
To address this limitation, we introduce Perfect Alternation (PA) as a reference coordination regime and propose six novel Alternation (ALT) metrics designed as temporally sensitive observables of coordination quality. Using Q-learning agents as a minimal adaptive diagnostic baseline, and comparing against random-policy null processes, we uncover a clear measurement failure: despite exhibiting deceptively high traditional metrics (e.g., reward fairness often exceeding $0.9$), learned policies perform up to $81\%$ below random baselines under ALT-variant evaluation—a deficit already present in the two-agent case and intensifying as $n$ grows.
@@ 1 Introduction
Coordination among multiple self-interested agents is a fundamental problem in complex systems, where local decision rules and asymmetric incentives give rise to emergent collective behavior Rankin et al. (2007); Axelrod (1984); Perolat et al. (2017). In the absence of centralized control, such systems exhibit interdependent feedback between individual actions and collective outcomes, producing diverse long-run coordination regimes Vainer and Kukacka (2021); Maruta and Okada (2012); Zhu et al. (2020); Shan et al. (2019); He et al. (2015); Tian et al. (2026); Pikovsky (2009). A central challenge is to characterize how efficient or equitable collective patterns emerge—and persist—purely through repeated interaction, adaptation, and feedback. Such coordination problems arise across diverse domains and turn-taking phenomena: decentralized access in wireless networks Shenoy (2007), conversational turn-taking in linguistics Sacks et al. (1974), animal communication and rhythmic alternation Takahashi et al. (2013); Greenfield et al. (2021); Ravignani et al. (2019); Pika et al. (2018), stochastic phenotype switching Rainey et al. (2011), vision-based robotic coordination Mezey et al. (2025), and emergent conversational norms in virtual environments Bonfert et al. (2022). Yet existing evaluation frameworks—including welfare measures from social choice theory and allocation criteria from fair division—remain largely insensitive to the temporal structure inherent in these coordination processes.
@@
Repeated interaction among egoistic agents can be naturally viewed as a discrete-time, stochastic, and history-dependent dynamical system Pikovsky (2009); Axelrod (1984). Early computational studies, most notably Axelrod’s investigation of cooperation among egoists without central authority Axelrod (1984), demonstrated that simple adaptive rules can generate stable behavioral regularities through repetition alone. From a game-theoretic perspective, learning rules induce adaptive trajectories in a high-dimensional strategy space, where long-run behavior corresponds to recurrent coordination regimes rather than one-shot equilibria Pikovsky (2009). Understanding which coordination patterns emerge, and how to distinguish them quantitatively, remains an open problem.
@@
The Battle of the Exes (BoE) game provides a minimal yet expressive model for studying such coordination dynamics Hawkins and Goldstone (2016). Unlike cooperative dilemmas, BoE is a coordination game with asymmetric payoffs, where agents must resolve conflict over preferred outcomes. In repeated play, the socially optimal solution is not static cooperation but temporal alternation, in which agents take turns accessing the higher-reward outcome Hawkins and Goldstone (2016); Papadopoulos and Sanchez-Fibla (2021); Papadopoulos (2020). This alternation corresponds to a structured temporal pattern—analogous to a periodic orbit—rather than a simple payoff equilibrium. As such, BoE offers a natural testbed for examining the emergence of turn-taking conventions as collective dynamical patterns.
@@
Despite extensive prior work on the repeated BoE, existing studies are almost exclusively restricted to the two-agent case Hawkins and Goldstone (2016); Puig Camps (2018); Gasparrini and Sanchez-Fibla (2018); Freire et al. (2020, 2023); Lopez Hernandez (2022). This limitation obscures coordination phenomena that arise in more realistic multi-agent settings, where competition among several agents induces richer and higher-dimensional dynamics. Moreover, considering BoE-derived multi-agent variants exposes a critical methodological gap: commonly used fairness and efficiency metrics are typically time-averaged and payoff-based Gini (1912); Theil (1967); Allison (1978); Jiang and Lu (2019); Perolat et al. (2017), rendering them insensitive to temporal ordering. As a result, such metrics cannot distinguish genuine turn-taking from monopolistic, random, or weakly correlated access patterns, particularly in multi-agent scenarios where high aggregate efficiency may coexist with poor coordination.
@@
In this study, we address these limitations by formalizing a BoE-derived multi-agent dynamic Battle of the Exes (MBoE) as a Markov game Littman (1994). To characterize turn-taking as a temporal phenomenon, we introduce the concept of a Perfect Alternation (PA) reference coordination regime for idealized turn-taking. Building on this benchmark, we propose six novel Alternation (ALT) metrics designed as time-series observables that explicitly capture temporal structure and coordination quality Papadopoulos and Sanchez-Fibla (2021); Papadopoulos et al. (2025). Through extensive numerical simulations with Q-learning agents—interpreted here as a discrete stochastic adaptive process and used as a minimal learning baseline—we demonstrate that traditional metrics can mask coordination failure, while ALT metrics reliably distinguish qualitatively different long-run regimes. Our results highlight the importance of temporally sensitive observables for analyzing coordination dynamics in strategic multi-agent interaction.
@@ 1.1 Main Contributions
    Perfect Alternation (PA): We introduce the concept of Perfect Alternation (PA) as a reference coordination regime for turn-taking tasks in multi-agent systems Papadopoulos (2020); Papadopoulos and Sanchez-Fibla (2021).
@@
    Novel Alternation (ALT) metrics: We propose six ALT measures that overcome the temporal insensitivity and coordination blindness of traditional fairness and efficiency metrics—their inherent inability to distinguish structured alternation from monopolistic or random access patterns, even when reporting values approaching 1.0. This dual limitation becomes severe and misleading as the number of agents grows, with traditional metrics masking severe coordination failure Papadopoulos and Sanchez-Fibla (2021); Papadopoulos et al. (2025).
@@
    Random baseline as null process: We establish random policy baselines as a statistical null hypothesis for coordination, demonstrating that high fairness or efficiency values can arise purely by chance—a methodological practice not consistently reported in prior BoE and closely related coordination studies Hawkins and Goldstone (2016); Puig Camps (2018); Gasparrini and Sanchez-Fibla (2018); Freire et al. (2020, 2023).
@@ 2.1 The Battle of the Exes Game
Hawkins and Goldstone investigated both ballistic and dynamic versions of BoE using human participants. Their experiments demonstrated that the dynamic formulation substantially increases the emergence of turn-taking behavior, with some pairs converging to near-perfect alternation, highlighting the role of spatiotemporal interaction in coordination dynamics Hawkins and Goldstone (2016).
@@ 2.2 Learning-Based Studies of BoE
Following these findings, several studies examined whether artificial agents could reproduce the alternation behavior observed in humans. Puig Camps (2018) investigated repeated BoE using independent Q-learning agents and reported convergence to stable deterministic strategies under specific learning-rate and exploration settings. Gasparrini and Sanchez-Fibla (2018) incorporated loss aversion into agents’ utility functions, accelerating convergence in the ballistic formulation and enabling alternation in the dynamic variant. Related modeling work explored control-based or episodic-memory agents in BoE-like settings, focusing on matching human efficiency/fairness/stability profiles rather than multi-agent scaling Freire et al. (2020, 2023); Lopez Hernandez (2022).
@@
Despite these advances, learning-based BoE studies remain almost exclusively restricted to the two-agent setting, where alternation reduces to binary role switching and random policies can achieve near-perfect Fairness, Reward Fairness, and TT Fairness, and high efficiency by chance alone. Evaluation relies primarily on time-averaged, payoff-based metrics such as efficiency and fairness, which suffer from temporal insensitivity and pattern ambiguity (random access can yield values indistinguishable from coordination) even in dyadic settings, with additional limitations—coordination blindness and exclusion of intermediate agents—emerging as $n$ grows. Moreover, random-policy baselines are not consistently reported as explicit null processes Hawkins and Goldstone (2016); Puig Camps (2018); Gasparrini and Sanchez-Fibla (2018); Freire et al. (2020, 2023), making it difficult to assess whether observed coordination reliably exceeds chance-level behavior. An exception is recent MSc thesis work that includes randomized control comparisons in dyadic settings Lopez Hernandez (2022); however, this comparison served primarily to validate model performance against human data rather than to establish a formal null hypothesis for multi-agent coordination assessment, leaving open the question of how severely traditional metrics can obscure coordination failure. Our study addresses these gaps by (i) formalizing random policies as an explicit statistical null process with quantitative Coordination Score and Relative Change methodology, (ii) scaling baselines to multi-agent settings ($n\in\{2,3,5,8,10\}$) across four experimental configurations, four outcome-based metrics, and six ALT variants, and (iii) demonstrating that traditional metrics can appear deceptively high even under random play—a finding that motivates the introduction of temporally sensitive ALT metrics. Building upon and substantially extending our early conceptual work on alternation-sensitive metrics and periodicity-based analysis Papadopoulos (2020); Papadopoulos and Sanchez-Fibla (2021); Papadopoulos et al. (2025), this paper introduces a rigorous theoretical framing, calibrated random baselines, and scaled multi-agent evaluation with refined PA and metric definitions.
@@ 2.4 Battle of the Exes as a Minimal Multi-Agent MARL Testbed
This BoE-derived formulation preserves the core incentive structure of BoE—mutual interference with asymmetric rewards—while providing a plausible multi-agent variant. The optimal collective behavior is no longer simple avoidance but periodic turn-taking across all agents. Whether and how such behavior emerges under decentralized learning—and how it should be evaluated—are the central questions addressed in the remainder of this manuscript.
@@ 2.5 Multi-Agent Battle of the Exes as a Markov Game
We now formally define the Multi-Agent Battle of the Exes (MBoE) as an episodic Markov game Littman (1994). This BoE-derived formulation constitutes a core contribution of this work, providing a plausible multi-agent variant of the classical two-agent coordination tension—mutual interference with asymmetric payoffs.
@@ 2.5.5 Reward Function
If all agents reach the terminal position simultaneously ($k=n$), no reward is assigned. This reward structure is inspired by the canonical two-agent Battle of the Exes, preserving its core coordination tension: agents benefit from exclusive access but suffer when competing for the same outcome. While the canonical two-agent BoE is defined over asymmetric destination preferences, the present formulation abstracts this structure into a single contested terminal with exclusivity-based rewards. This abstraction preserves the essential coordination tension—mutual interference under asymmetric benefit—while enabling systematic scaling to the multi-agent regime.
@@ 3 Traditional Multi-Agent Fairness Measures
Multi-agent fairness research distinguishes between outcome fairness (cumulative payoff distribution) and procedural fairness (temporal distribution of rewards). Common outcome measures include the Gini coefficient Gini (1912), Theil index Theil (1967), and Coefficient of Variation Allison (1978), while recent multi-agent work often uses payoff-based fairness or equality variants Jiang and Lu (2019); Perolat et al. (2017). In the fair division literature, allocation mechanisms are evaluated by properties such as envy-freeness, proportionality, and strategyproofness Suksompong and Teh (2023); Camacho et al. (2023), yet these criteria remain distributional and do not account for the temporal sequence of allocations. Similarly, axiomatic approaches to cooperative surplus sharing emphasize distributional fairness and consistency Calleja and Llerena (2020), but do not address the temporal ordering of payoffs. Consequently, these metrics fail to capture when agents receive rewards—critical for turn-taking tasks where an agent monopolizing early episodes violates fairness even if cumulative payoffs equalize. These shortcomings already appear in the two-agent BoE setting; see Supplementary Section S3 for extended discussion. This limitation motivates the alternation-sensitive measures developed in Section 4.
@@ 3.1 Limitations of Traditional Metrics in BoE and MBoE
Defined as $F_{\nu}^{TT}=\min_{i}t_{i}/\max_{i}t_{i}$ Hawkins and Goldstone (2016). Key limitations: (1) still ignores intermediate agents, and (2) high $t_{i}$ can reflect frequent collisions rather than genuine alternation.
@@
Defined as $F_{\nu}^{R}=\min_{i}p_{i}/\max_{i}p_{i}$ Hawkins and Goldstone (2016); Puig Camps (2018); Gasparrini and Sanchez-Fibla (2018), comparing cumulative payoffs. Key limitations: (1) high values can result from tie rewards without turn-taking, and (2) a dominant subgroup can still yield high Reward Fairness.
@@
Combining Efficiency with Fairness or TT Fairness does not even guarantee a fair outcome interpretation: Fairness ignores intermediate agents, while TT Fairness conflates exclusive wins with collision-induced arrivals. Only Reward Fairness, when combined with high Efficiency, can indicate equitable payoff distribution—but even then, temporal structure remains invisible. This double inadequacy—failing both as fairness indicators and as coordination measures—motivates the development of ALT metrics. Table 2 illustrates that interpretable patterns emerge only at extremes; at intermediate values—common in practice—these metrics fail to distinguish coordination quality. Random-policy baselines are not consistently reported as explicit null references in prior BoE-related studies Hawkins and Goldstone (2016); Puig Camps (2018); Gasparrini and Sanchez-Fibla (2018); Freire et al. (2020, 2023); we therefore include them as an explicit null reference for coordination assessment.

### 10 Mar 2026 v2

@@ Abstract
Multi-agent coordination dilemmas expose a fundamental tension between individual optimization and collective welfare, yet characterizing such coordination requires metrics sensitive to temporal structure and collective dynamics. As a diagnostic testbed, we study a BoE-derived multi-agent variant of the Battle of the Exes, formalizing it as a Markov game in which turn-taking emerges as a periodic coordination regime. Conventional outcome-based metrics (e.g., efficiency and min/max fairness) are temporally blind (they cannot distinguish structured alternation from monopolistic or random access patterns) and fairness ratios lose discriminative power as $n$ grows, obscuring inequities.
To address this limitation, we introduce Perfect Alternation (PA) as a reference coordination regime and propose six novel Alternation (ALT) metrics designed as temporally sensitive observables of coordination quality. Using Q-learning agents as a minimal adaptive diagnostic baseline, and comparing against random-policy null processes, we uncover a clear measurement failure: despite exhibiting deceptively high traditional metrics (e.g., reward fairness often exceeding $0.9$), learned policies perform up to $81\%$ below random baselines under ALT-variant evaluation, a deficit already present in the two-agent case and intensifying as $n$ grows.
@@ 1 Introduction
Coordination among multiple self-interested agents is a fundamental problem in complex systems, where local decision rules and asymmetric incentives give rise to emergent collective behavior (Rankin et al., 2007; Axelrod, 1984; Perolat et al., 2017). In the absence of centralized control, such systems exhibit interdependent feedback between individual actions and collective outcomes, producing diverse long-run coordination regimes (Vainer and Kukacka, 2021; Maruta and Okada, 2012; Zhu et al., 2020; Shan et al., 2019; He et al., 2015; Tian et al., 2026; Pikovsky, 2009). A central challenge is to characterize how efficient or equitable collective patterns emerge (and persist) purely through repeated interaction, adaptation, and feedback. Such coordination problems arise across diverse domains and turn-taking phenomena: decentralized access in wireless networks (Shenoy, 2007), conversational turn-taking in linguistics (Sacks et al., 1974), animal communication and rhythmic alternation (Takahashi et al., 2013; Greenfield et al., 2021; Ravignani et al., 2019; Pika et al., 2018), stochastic phenotype switching (Rainey et al., 2011), vision-based robotic coordination (Mezey et al., 2025), and emergent conversational norms in virtual environments (Bonfert et al., 2022). Yet existing evaluation frameworks (including welfare measures from social choice theory and allocation criteria from fair division) remain largely insensitive to the temporal structure inherent in these coordination processes.
@@
Repeated interaction among egoistic agents can be naturally viewed as a discrete-time, stochastic, and history-dependent dynamical system (Pikovsky, 2009; Axelrod, 1984). Early computational studies, most notably Axelrod’s investigation of cooperation among egoists without central authority (Axelrod, 1984), demonstrated that simple adaptive rules can generate stable behavioral regularities through repetition alone. From a game-theoretic perspective, learning rules induce adaptive trajectories in a high-dimensional strategy space, where long-run behavior corresponds to recurrent coordination regimes rather than one-shot equilibria (Pikovsky, 2009). Understanding which coordination patterns emerge, and how to distinguish them quantitatively, remains an open problem.
@@
The Battle of the Exes (BoE) game provides a minimal yet expressive model for studying such coordination dynamics (Hawkins and Goldstone, 2016). Unlike cooperative dilemmas, BoE is a coordination game with asymmetric payoffs, where agents must resolve conflict over preferred outcomes. In repeated play, the socially optimal solution is not static cooperation but temporal alternation, in which agents take turns accessing the higher-reward outcome (Hawkins and Goldstone, 2016; Papadopoulos and Sanchez-Fibla, 2021; Papadopoulos, 2020). This alternation corresponds to a structured temporal pattern (analogous to a periodic orbit) rather than a simple payoff equilibrium. As such, BoE offers a natural testbed for examining the emergence of turn-taking conventions as collective dynamical patterns.
@@
Despite extensive prior work on the repeated BoE, existing studies are almost exclusively restricted to the two-agent case (Hawkins and Goldstone, 2016; Puig Camps, 2018; Gasparrini and Sanchez-Fibla, 2018; Freire et al., 2020, 2023; Lopez Hernandez, 2022). This limitation obscures coordination phenomena that arise in more realistic multi-agent settings, where competition among several agents induces richer and higher-dimensional dynamics. Moreover, considering BoE-derived multi-agent variants exposes a critical methodological gap: commonly used fairness and efficiency metrics are typically time-averaged and payoff-based (Gini, 1912; Theil, 1967; Allison, 1978; Jiang and Lu, 2019; Perolat et al., 2017), rendering them insensitive to temporal ordering. As a result, such metrics cannot distinguish genuine turn-taking from monopolistic, random, or weakly correlated access patterns, particularly in multi-agent scenarios where high aggregate efficiency may coexist with poor coordination.
@@
In this study, we address these limitations by formalizing a BoE-derived multi-agent dynamic Battle of the Exes (MBoE) as a Markov game (Littman, 1994). To characterize turn-taking as a temporal phenomenon, we introduce the concept of a Perfect Alternation (PA) reference coordination regime for idealized turn-taking. Building on this benchmark, we propose six novel Alternation (ALT) metrics designed as time-series observables that explicitly capture temporal structure and coordination quality (Papadopoulos and Sanchez-Fibla, 2021; Papadopoulos et al., 2025). Through extensive numerical simulations with Q-learning agents (interpreted here as a discrete stochastic adaptive process and used as a minimal learning baseline), we demonstrate that traditional metrics can mask coordination failure, while ALT metrics reliably distinguish qualitatively different long-run regimes. Our results highlight the importance of temporally sensitive observables for analyzing coordination dynamics in strategic multi-agent interaction.
@@ 1.1 Main Contributions
    Perfect Alternation (PA): We introduce the concept of Perfect Alternation (PA) as a reference coordination regime for turn-taking tasks in multi-agent systems (Papadopoulos, 2020; Papadopoulos and Sanchez-Fibla, 2021).
@@
    Novel Alternation (ALT) metrics: We propose six ALT measures that overcome the temporal insensitivity and coordination blindness of traditional fairness and efficiency metrics (their inherent inability to distinguish structured alternation from monopolistic or random access patterns, even when reporting values approaching 1.0). This dual limitation becomes severe and misleading as the number of agents grows, with traditional metrics masking severe coordination failure (Papadopoulos and Sanchez-Fibla, 2021; Papadopoulos et al., 2025).
@@
    Random baseline as null process: We establish random policy baselines as a statistical null hypothesis for coordination, demonstrating that high fairness or efficiency values can arise purely by chance, a methodological practice not consistently reported in prior BoE and closely related coordination studies (Hawkins and Goldstone, 2016; Puig Camps, 2018; Gasparrini and Sanchez-Fibla, 2018; Freire et al., 2020, 2023).
@@ 2.1 The Battle of the Exes Game
Hawkins and Goldstone investigated both ballistic and dynamic versions of BoE using human participants. Their experiments demonstrated that the dynamic formulation substantially increases the emergence of turn-taking behavior, with some pairs converging to near-perfect alternation, highlighting the role of spatiotemporal interaction in coordination dynamics (Hawkins and Goldstone, 2016).
@@ 2.2 Learning-Based Studies of BoE
Following these findings, several studies examined whether artificial agents could reproduce the alternation behavior observed in humans. Puig Camps (2018) investigated repeated BoE using independent Q-learning agents and reported convergence to stable deterministic strategies under specific learning-rate and exploration settings. Gasparrini and Sanchez-Fibla (2018) incorporated loss aversion into agents’ utility functions, accelerating convergence in the ballistic formulation and enabling alternation in the dynamic variant. Related modeling work explored control-based or episodic-memory agents in BoE-like settings, focusing on matching human efficiency/fairness/stability profiles rather than multi-agent scaling (Freire et al., 2020, 2023; Lopez Hernandez, 2022).
@@
Despite these advances, learning-based BoE studies remain almost exclusively restricted to the two-agent setting, where alternation reduces to binary role switching and random policies can achieve near-perfect Fairness, Reward Fairness, and TT Fairness, and high efficiency by chance alone. Evaluation relies primarily on time-averaged, payoff-based metrics such as efficiency and fairness, which suffer from temporal insensitivity and pattern ambiguity (random access can yield values indistinguishable from coordination) even in dyadic settings, with additional limitations (coordination blindness and exclusion of intermediate agents) emerging as $n$ grows. Moreover, random-policy baselines are not consistently reported as explicit null processes (Hawkins and Goldstone, 2016; Puig Camps, 2018; Gasparrini and Sanchez-Fibla, 2018; Freire et al., 2020, 2023), making it difficult to assess whether observed coordination reliably exceeds chance-level behavior. An exception is recent MSc thesis work that includes randomized control comparisons in dyadic settings (Lopez Hernandez, 2022); however, this comparison served primarily to validate model performance against human data rather than to establish a formal null hypothesis for multi-agent coordination assessment, leaving open the question of how severely traditional metrics can obscure coordination failure. Our study addresses these gaps by (i) formalizing random policies as an explicit statistical null process with quantitative Coordination Score and Relative Change methodology, (ii) scaling baselines to multi-agent settings ($n\in\{2,3,5,8,10\}$) across four experimental configurations, four outcome-based metrics, and six ALT variants, and (iii) demonstrating that traditional metrics can appear deceptively high even under random play—a finding that motivates the introduction of temporally sensitive ALT metrics. Building upon and substantially extending our early conceptual work on alternation-sensitive metrics and periodicity-based analysis (Papadopoulos, 2020; Papadopoulos and Sanchez-Fibla, 2021; Papadopoulos et al., 2025), this paper introduces a rigorous theoretical framing, calibrated random baselines, and scaled multi-agent evaluation with refined PA and metric definitions.
@@ 2.4 Battle of the Exes as a Minimal Multi-Agent MARL Testbed
This BoE-derived formulation preserves the core incentive structure of BoE (mutual interference with asymmetric rewards) while providing a plausible multi-agent variant. The optimal collective behavior is no longer simple avoidance but periodic turn-taking across all agents. Whether and how such behavior emerges under decentralized learning (and how it should be evaluated) are the central questions addressed in the remainder of this manuscript.
@@ 2.5 Multi-Agent Battle of the Exes as a Markov Game
We now formally define the Multi-Agent Battle of the Exes (MBoE) as an episodic Markov game (Littman, 1994). This BoE-derived formulation constitutes a core contribution of this work, providing a plausible multi-agent variant of the classical two-agent coordination tension (mutual interference with asymmetric payoffs).
@@ 2.5.5 Reward Function
If all agents reach the terminal position simultaneously ($k=n$), no reward is assigned. This reward structure is inspired by the canonical two-agent Battle of the Exes, preserving its core coordination tension: agents benefit from exclusive access but suffer when competing for the same outcome. While the canonical two-agent BoE is defined over asymmetric destination preferences, the present formulation abstracts this structure into a single contested terminal with exclusivity-based rewards. This abstraction preserves the essential coordination tension (mutual interference under asymmetric benefit) while enabling systematic scaling to the multi-agent regime.
@@ 3 Traditional Multi-Agent Fairness Measures
Multi-agent fairness research distinguishes between outcome fairness (cumulative payoff distribution) and procedural fairness (temporal distribution of rewards). Common outcome measures include the Gini coefficient (Gini, 1912), Theil index (Theil, 1967), and Coefficient of Variation (Allison, 1978), while recent multi-agent work often uses payoff-based fairness or equality variants (Jiang and Lu, 2019; Perolat et al., 2017). In the fair division literature, allocation mechanisms are evaluated by properties such as envy-freeness, proportionality, and strategyproofness (Suksompong and Teh, 2023; Camacho et al., 2023), yet these criteria remain distributional and do not account for the temporal sequence of allocations. Similarly, axiomatic approaches to cooperative surplus sharing emphasize distributional fairness and consistency (Calleja and Llerena, 2020), but do not address the temporal ordering of payoffs. Consequently, these metrics fail to capture when agents receive rewards, critical for turn-taking tasks where an agent monopolizing early episodes violates fairness even if cumulative payoffs equalize. These shortcomings already appear in the two-agent BoE setting; see Supplementary Section S3 for extended discussion. This limitation motivates the alternation-sensitive measures developed in Section 4.
@@ 3.1 Limitations of Traditional Metrics in BoE and MBoE
Defined as $F_{\nu}^{TT}=\min_{i}t_{i}/\max_{i}t_{i}$ (Hawkins and Goldstone, 2016). Key limitations: (1) still ignores intermediate agents, and (2) high $t_{i}$ can reflect frequent collisions rather than genuine alternation.
@@
Defined as $F_{\nu}^{R}=\min_{i}p_{i}/\max_{i}p_{i}$ (Hawkins and Goldstone, 2016; Puig Camps, 2018; Gasparrini and Sanchez-Fibla, 2018), comparing cumulative payoffs. Key limitations: (1) high values can result from tie rewards without turn-taking, and (2) a dominant subgroup can still yield high Reward Fairness.
@@
Combining Efficiency with Fairness or TT Fairness does not even guarantee a fair outcome interpretation: Fairness ignores intermediate agents, while TT Fairness conflates exclusive wins with collision-induced arrivals. Only Reward Fairness, when combined with high Efficiency, can indicate equitable payoff distribution—but even then, temporal structure remains invisible. This double inadequacy (failing both as fairness indicators and as coordination measures) motivates the development of ALT metrics. Table 2 illustrates that interpretable patterns emerge only at extremes; at intermediate values—common in practice—these metrics fail to distinguish coordination quality. Random-policy baselines are not consistently reported as explicit null references in prior BoE-related studies (Hawkins and Goldstone, 2016; Puig Camps, 2018; Gasparrini and Sanchez-Fibla, 2018; Freire et al., 2020, 2023); we therefore include them as an explicit null reference for coordination assessment.